(5,4) (6,5) (7,6) (8.7) (9,8)

Consider the five-data pair (x, y) and

- Find the first and second derivatives exactly at x = c. (c is any x in your data!)
Answer #1

Given 07 6 os 9 f(x) 4 5. 6 . 7 8 Xith Suppose we want to find first derivative at x= xi then f (xi)= f (x+1) - f (x,;-) xilNow Suppose we want to find and derivative at x= x then f (xi) = f (tax)_2 f(xi) + f(x2-x) (4x)2 Ax=1 an given Question [ Ax